Two points A and B move from rest along a straight line with constant acceleration f and f' respectively. If A takes m sec. more than B and describes ‘n’ units more than B in acquiring the same speed then ?

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Given two points A and B move from rest along a straight line with constant acceleration f and f' respectively. If A takes m sec. more than B and describes ‘n’ units more than B in acquiring the same speed then
v2 = 2f(d + n) = 2f′d
v = f′(t) = (m + t)f
eliminate d and m we get
(f′ - f)n = (1/2) ff’m2
Thanks and Regards
Shaik Aasif
askIITians faculty